40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 154 of 330  Not logged in ELOG logo
ID Date Author Type Category Subject
  8888   Mon Jul 22 06:58:17 2013 LisaSummaryloreAngel of the Y End Table?

Quote:

 Trying to take an image or movie of the ETMY Transmon cam, we got instead this attached image.

I think it is just some scattered green light, but others in the control room think that it is a message from somewhere or someone...

 It is not an angel, it is clearly a four leaf clover (also known as "quadrifoglio"). It is very rare, it brings good luck!

Attachment 1: image.jpg
image.jpg
  8887   Mon Jul 22 03:10:41 2013 ranaSummaryloreAngel of the Y End Table?

 Trying to take an image or movie of the ETMY Transmon cam, we got instead this attached image.

I think it is just some scattered green light, but others in the control room think that it is a message from somewhere or someone...

Attachment 1: asdasd.jpg
asdasd.jpg
  8886   Mon Jul 22 03:09:51 2013 AnnalisaUpdateGreen LockingY Arm cavity scan

Yesterday and today I was in the lab doing many cavity scan.

First I did many measurement with the cavity aligned in order to get the position of the 00 modes, then I misaligned the beam in many different ways to enhance the higher order modes.

In particular, I first misaligned the mode cleaner to make the beam clipping into the Faraday. To do this, I set to 0 the WFS gain, but I left the autolocker still enabled. In this way, the autolocker couldn't bring the mirrors back to the aligned position.

Then I misaligned also the TT2 to get even more HOMs.

Eventually, Rana came and we misaligned TT1 to clip the beam, and using TT2 we aligned back the beam to the arm.

To increase the SNR, we changed the gain of the TRY PD, setting it to 20dB (which corresponds to a factor 100 in digital scale)

I attached one scan that I did with Rana on Sunday night. I could not upload a better resolution image because the file size was too big, but here's the path to find all of the scans:

../users/annalisa/sweep/Yarm

There are many folders, one per each day I measured. In each folder there are measurements relative to aligned cavity, Pitch and Yaw misalignment.

 RXA EDIT:

The PDA520 used for TRY was set to 0 dB analog gain. This corresponds to ~500 counts out of 32768. The change to 20 dB actually increases the gain by 100. This makes the single arm lock saturate at ~25000 counts (obviously in analog before the ADC). The right setting for our usual running is probably 10 dB.

For the IMC WFS, we had disabled the turn on in the autolocker to use the IMC to steer the beam in the FI, but that was a flop (not enough range, not enough lever arm). In the end, I think we didn't get any clipping.

 

 

Attachment 1: Screenshot-scan3_0722.pdf_-_Adobe_Reader.png
Screenshot-scan3_0722.pdf_-_Adobe_Reader.png
  8885   Fri Jul 19 23:35:52 2013 KojiSummaryLSCTRY DC locking

The Y arm was locked with the TRY DC signal.

The handing off process is too complicated because there is no path from ALS to the LSC error.


 The TRY DC error signal & the gain determination

- The error signal was produced by the operation 1/SQRT(TRY) - OFFSET. The initial offset was -5.

- The sign of the TRY DC error signal depends on which side of the resonance the arm is.
  By looking at the strip chart, I determined that the sign is opposite of the ALS.
  The ALS had the gain of -25, so the TRY control gain was to be positive.

- From the strip chart on the previous entry , the slope difference between the PDH error and the TRY DC error was x500.
  The arm control with POY11 PDH had the gain of 0.2. So the target gain for the TRY DC was determined to be +100.

Handing off

- The arm was stabilized by ALS. The ALS gain was -25 with FM2/3/5/6/7/10

- YARM configuration: no trigger / no FM trigger / gain =+0 / FM5 ON / OFFSET -5

- Start handing off:
  YARM: Turned up the gain to +50

- ALS: Turned off FM6/7

- YARM: Turned on FM6/7

- ALS: Turned off FM2

- YARM: Turned on FM4

- ALS: Turned off FM3/10

- YARM: Turned on FM2/3/8/9 ON

- ALS: Reduced the gain to -15

- YARM: Increased the gain to +70

- ALS: Reduced the gain to 0

- YARM: Increased the gain to +100

HANDING OFF - DONE

Changing the offset

The offset of -5 gave the TRY of <0.1.

The detuning was reduced by giving the offset of -4. TRY went up to ~.1

The offset of -3 made TRY 0.13

The offset of -2 made TRY 0.25

The offset of -1.5 made TRY 0.4. And the arm could not be held by this error signal anymore.

 

 

Attachment 1: TRY_locking.png
TRY_locking.png
  8884   Fri Jul 19 23:35:31 2013 gautamConfigurationendtable upgradePreliminary Calibration of PZT

 I recalibrated the QPD today as I had shifted its position a little. I then identified the linear range of the QPD and performed a preliminary calibration of the Piezo tip-tilt within this range.

 


Details:

-I recalibrated the QPD as I had shifted it around a little in order to see if I could move it to a position such that I could get the full dynamic range of the piezo tilt within the linear regime of the QPD. This proved difficult because there are two reflections from the mirror (seeing as it is AR coated for 532nm and I am using a red laser). At a larger separation, these diverge and the stray spot does not bother me. But it does become a problem when I move the QPD closer to the mirror (in an effort to cut down the range in which the spot on the QPD moves). In any case, I had moved the QPD till it was practically touching the mirror, and even then, could not get the spot motion over the full range of the PZTs motion to stay within the QPD's linear regime (as verified by applying a 20Vpp 1Hz sine wave to the PZT driver board and looking at the X and Y outputs from the QPD amplifier. 

-So I reverted to a configuration in which the QPD was ~40cm away from the mirror (measured using a measuring tape).

-The new calibration constants are as follows (see attached plots):

X-Coordinate: -3.43 V/mm
Y-Coordinate: -3.41 V/mm


-I then determined the linear range of the QPD to be when the output was in the range [-0.5V 0.5V]. 

-Next, at Jenne's suggestion, I decided to do a preliminary calibration of the PZT within this linear range. I used an SR function generator to supply an input voltage to the PZT driver board's input (connected to Channel 1 of the piezo). In order to supply a DC voltage, I set a DC offset, and set the signal amplitude to 0V. I then noted the X and Y-coordinate outputs, being sure to run through the input voltages in a cyclic fashion as one would expect some hysteresis. 

-I did this for both the pitch and yaw inputs, but have only superficially analysed the latter case (I will put up results for the former later). 

 


Comments:

-There is indeed some hysteresis, though the tilt seems to vary linearly with the input voltage. I have not yet included a calibration constant as I wish to perform this calibration over the entire dynamic range of the PZT. 

 

-There is some residual coupling between the pitch and yaw motion of the tip tilt, possibly due to its imperfect orientation in the holder (I have yet to account for the QPD's tilt).

 

-I have not included a graphical representation here, but there is significantly more pitch to yaw coupling when my input signal is applied to the tip-tilts pitch input (Channel 2), as compared to when it is input to channel 1. It is not clear to me why this is so.

 

-I have to think of some smart way of calibrating the PZT over its entire range of motion, keeping the spot in the QPD's linear regime throughout. One idea is to start at one extreme (say with input voltage -10V), and then perform the calibration, re-centering the spot to 0 on the QPD each time the QPD amp output reaches the end of its linear regime. I am not sure if this will work, but it is worth a shot. The other option is to replace the red laser with a green laser (from one of the laser pointers) in the hope that multiple reflections will be avoided from the mirror. Then I will have to recalibrate the set up, and see if I can get the QPD close enough to the mirror such that the spot stays within the linear regime of the QPD. More investigation needs to be done.


Plots:

QPD Calibration Plots:
QPD-XCalib.pdf            QPD-YCalib.pdf

 

 

Piezo tilt vs input voltage plots:

 

                                                          Yaw Tilt                                                                                                                                         Pitch Tilt

Piezo_Yaw_Calib-in_QPD_linear_range-Yaw_tilt.pdf               Piezo_Yaw_Calib-in_QPD_linear_range-Pitch_tilt.pdf 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

  8883   Fri Jul 19 22:51:40 2013 gautamConfigurationendtable upgradeSecond mirror glued to PZT and mounted

 

I mounted the second PZT in a modified mount, and then glued a 1-inch Y2 mirror on it using superglue.

 Details:

-The mirror is a Laseroptik 1-inch, Y2 mirror with HR and AR coatings for 532 nm light.

-The procedure for mounting the mirror was the same as detailed in elog 8874. This time, I tried to orient the Piezo such that the four screws on the back face coincided with the horizontal and vertical axes, as this appeared to (somewhat) decouple the pitch and yaw motion of the tip-tilt on the first PZT.

-One thing I forgot to mention in the earlier elog: it is best to assemble the mount fully before inserting the tip-tilt into it and gluing the mirror to the tip-tilt. In particular, the stand should be screwed onto the mount before inserting the tip-tilt into the holder, as once it is in, it will block the hole through which one can screw the stand onto the mount. 

-I have placed the mirror on the SP table along with the rest of my QPD/Piezo calibration setup. I will attempt to calibrate this second PZT once I am done with the first one. 
 
Here is an image of the assembly:
Piezo-mirror_2.JPG

 

 

 

 

 

 

 

  8882   Fri Jul 19 22:35:06 2013 KojiSummaryLSCVarious Arm signal (Yarm)

The StripTool plot attached below shows various arm signals measured with the Y arm cavity swept using ALS.

Yellow: TRY

Blue: ALS additive OFFSET to the error signal

Red: Raw PDH error signal (POY11I)

Purple: Linearized PDH error (POY11/TRY)

Green: 1/Sqrt(TRY)-5 (No normalization)

Inverse Sqrt of the TRY had been implemented when this LSC controller was first coded.
It is confirmed that the calculation is working correctly.

Attachment 1: various_arm_signal.png
various_arm_signal.png
  8881   Fri Jul 19 14:04:24 2013 KojiUpdateCDSCDS FE not happy

daqd was restarted.


- tried telnet fb 8088 on rossa => same error as manasa had

- tried telnet fb 8087 on rossa => same result

- sshed into fb ssh fb

- tried to find daqpd by ps -def | grep daqd => not found

- looked at wiki https://wiki-40m.ligo.caltech.edu/New_Computer_Restart_Procedures?highlight=%28daqd%29

- the wiki page suggested the following command to run daqd /opt/rtcds/caltech/c1/target/fb/daqd -c ./daqdrc &

- ran ps -def | grep nds => already exist. Left untouched.

- Left fb.

- tried telnet fb 8087 on rossa => now it works

  8880   Fri Jul 19 12:23:34 2013 manasaUpdateCDSCDS FE not happy

I found CDS rt processes in red. I did 'mxstreamrestart' from the medm. It did not help. Also ssh'd into c1iscex and tried 'mxstreamrestart' from the command line. It did not work either.

I thought restarting frame builder would help. I ssh'd to fb. But when I try to restart fb I get the following error:

controls@fb ~ 0$ telnet fb 8088
Trying 192.168.113.202...
telnet: connect to address
192.168.113.202: Connection refused

 

Screenshot-Untitled_Window.png

  8879   Fri Jul 19 12:02:18 2013 manasaUpdateLSCPRMI+Y arm ALS success!

Data retrieved using getdata (30 minutes trend) saved at

/users/manasa/data/130717/PRMI_YALS

  8878   Fri Jul 19 12:00:12 2013 manasaUpdateGreen LockingALS Y performance with the new whitening filter

Quote:


(Manasa downloaded the 2k sampled data so that we can use this for presentations.)

 Path to data (retreived using getdata)

/users/manasa/data/130717/YALS_scan

  8877   Thu Jul 18 23:34:40 2013 gautamConfigurationendtable upgradeCoarse adjustment of PZT axes orientation in mount

 I have managed to orient the PZT in the mount such that its axes are approximately aligned with the vertical and the horizontal. 

In the process, I discovered that the 4 screws on the back face of the PZT correspond to the location of the piezoelectric stacks beneath the tip-tilt platform. The PZT can therefore be oriented during the mounting process itself, before the mirror is glued onto the tip-tilt platform.

In order to verify that the pitch and yaw motion of the mirror have indeed been roughly decoupled, I centred the spot on the QPD, fed to the 'pitch' input of the PZT driver board (connected to channel 1 of the PZT) a 10 Vpp, 1 Hz sine wave from the SR function generator (having turned all the other relevant electronics, HV power supply etc ON. The oscilloscope trace of the output observed on the QPD is shown. The residual fluctuation in the Y-coordinate (blue trace) is I believe due to the tilt in the QPD, and also due to the fact that the PZT isnt perfectly oriented in the mount.

It looks like moving the tip-tilt through its full range of motion takes us outside the linear regime of the QPD calibration. I may have to rethink the calibration setup to keep the spot on the QPD in the linear range if the full range is to be calibrated, possibly decrease the distance between the mirror and the QPD. Also, in the current orientation, CH1 on the PZT controls YAW motion, while CH2 controls pitch.

Oscilloscope Trace:

Yellow: X-coordinate

Blue: Y-coordinate

PZT_out_waveform.jpg

  8876   Thu Jul 18 21:45:36 2013 CharlesUpdateISSISS - Full Schematic

 Here I have included the full schematic (so far) of the proposed ISS. There are two sheets: the first schematic details the filter stages and their accompanying circuitry while the second schematic details the RMS threshold detection and subsequent triggering.

The first schematic is fairly self explanatory as to what different portions do, and I have annotated much of the second schematic as there are some non-traditional components etc.

I have not yet included some mechanism to adjust the threshold voltage in real time or any of the power regulation, but these should follow fairly quickly.

Attachment 1: 40mServo_v1.pdf
40mServo_v1.pdf 40mServo_v1.pdf
  8875   Thu Jul 18 21:12:58 2013 gautamConfigurationendtable upgradeAI Board-D000186-All channels modified

 

 I carried some further modifications and tests to the AI Board. Details and observations here:

  • I switched out the resistors for all the remaining 7 channels, using the same substitutions as detailed here
  • I then verified that the modified transfer function for all 8 channels using the SR785. I did not collect data for all the channels as netgpib was taking ages, but I did use the cursor on the screen to verify the position of the first notch at ~64 kHz. I noticed that all the channels did not have the lowest point of the notch at the same frequency. Rather, (at least on the screen), this varied between 63kHz and 67kHz. I would put this down to component tolerance. Assuming 5% tolerance shifts the theoretical notch frequency from 66268 Hz to 63112 Hz. 
  • After verifying the transfer functions, I went to 1Y4 and plugged the AI board into the eurocrate. I then connected the input of the AI board to the DAC output using my custom ribbon cable. Next, I used the excitation points set up earlier to send a 1 kHz, 32000 counts amplitude sine wave through the channels one at a time. I monitored the output using an oscilloscope and the LEMO monitor channels on the front panel of the board.
  • I found that the single-ended output of the AI board swings between -10 V and 10 V (w.r.t ground, oscilloscope trace attached). This is good because this is the range of input voltage to the PZT driver boards required to realize the full actuation range of the PZTs.
  • I also verified that the connections on the custom ribbon cable are correct (channel map was right) and that there were no accidental shorts (I checked other channels' output monitor while driving one channel). 

I think the board is okay to be used now.

AI_Board_Output.jpg

 

  8874   Thu Jul 18 20:20:52 2013 gautamConfigurationendtable upgradeFirst mirror glued to PZT and mounted in modified mounts

 

 Yesterday, I mounted the first PZT in one of the modified mounts, and then glued a 2-inch Y2 mirror on it using superglue.

Details:

-The mirror is a 2-inch, Y2 mirror with HR and AR coatings for 532 nm light.

-The AR side of the mirror had someone's fingerprint on it, which I removed (under Manasa's guidance) using tweezers wrapped in lens cleaning paper, and methanol.

-Before gluing the mirror, I had to assemble the modified mount. Manasa handed over the remaining parts of the mounts (which are now in my newly acquired tupperware box along with all the other Piezo-related hardware). I took the one labelled A, and assembled the holder part. I then used one of the new mounts (2.5 inches, these are with the clean mounts in a cardboard box in the cupboard holding the green optics along the Y-arm) and mounted the holder on it. 

-Having assembled the mount, I inserted the piezo tip-tilt into the holder. The wedge that the machine shop supplied is useful (indeed required) for this. 

-I then cleaned the AR surface of the mirror and the top-surface of the tip-tilt. 

-The gluing was done using superglue which Steve got from the bookstore (the remaining tube is in the small fridge). We may glue the other mirror using epoxy. I placed 4 small drops of superglue on the tip-tilt's top surface, placed the mirror with its AR face in contact with the piezo, and applied some pressure for a short while until the glue spread out fairly evenly. I then left the whole setup to dry for about half an hour.

-Steve suggested using a reference piece (I used two small bolts) to verify when the glue had dried.

-Finally, I attached the whole assembly to a base.

Here it is in action in my calibration setup (note that it has not been oriented yet. i.e. the two perpendicular axes of the piezo are for the time being arbitrarily oriented. And maybe the spreading of the glue wasn't that even after all...):

Piezo-mirror.jpg

 

Sidenote:

Yesterday, while setting stuff up, I tested the piezo with a 0.05 Hz, 10Vpp input from the SR function generator just to see if it works, and also to verify that I had set up all my electronics correctly. Though the QPD was at this point calibrated, I did observe periodic motion of both the X and Y outputs of my QPD amp! Next step- calibration... 

  8873   Thu Jul 18 19:09:08 2013 gautamConfigurationendtable upgradeQPD Calibration for PZT Calibration

Summary 

I have been working on setting up a QPD which can eventually be used to calibrate the PZT, and also orient the PZT in the mount such that the pitch and yaw axes roughly coincide with the vertical and horizontal.

The calibration constants have been determined to be:

X-axis: -3.69 V/mm

Y-axis: -3.70V/mm

Methodology:

I initially tried using the QPD setup left behind by Chloe near MC2, but this turned out to be dysfunctional. On opening out the QPD, I found that the internal circuitry had some issues (shorts in the wrong places etc.) Fortunately, Steve was able to hand me another working unit. For future reference, there are a bunch of old QPDs which I assume are functional in the cabinet marked 'Old PDs' along the Y-arm. 

I then made a circuit with which to read out the X and Y coordinates from the QPD. This consists of 4 buffer amplifiers (one for each quadrant), and 3 summing amplifiers (outputs are A+B+C+D = sum, B+C-A-D = Y-coordinate, and A+B-C-D = X-coordinate) that take the appropriate linear combinations of the 4 quadrants to output a voltage that may be calibrated against displacement of the QPD. 

The output from the QPD is via a sub-D connector on the side of the pomona box enclosing the PD and the circuitry, with 7 pins- 3 for power lines, and 4 for the 4 quadrants of the QPD. It was a little tricky to figure the pin-out for this connector, as there was no way to use continuity checking to map the pins to quadrants. Therefore, I used a laser pointer, and some trial and error (i.e. shine the light on a given quadrant, and check the sign of the X and Y voltages on an oscilloscope) to map the pin outs. Steve tells me that these QPDs were made long before colour code standardisation, but I note here the pin outs in any case for future reference (the quadrant orientations are w.r.t the QPD held with all the circuitry above it, with the active surface facing me):

Red= +Vcc

Black= -Vcc

Green = GND

Blue = Upper Left Quadrant

White = Upper Right Quadrant

Purple = Lower Left Quadrant

Grey = Lower Right Quadrant

Chloe had noted that there was some issue with the voltage regulators on her circuit (overheating) but I suspect this may have been due to the faulty internal circuitry. Also, she had used 12 V regulators. I checked the datasheet of the QPD, Op-Amp LF347 (inside the pomona box) and the OP27s on my circuit, and found that they all had absolute maximum ratings above 18V, so I used 15V voltage regulators. The overheating problem was not a problem anymore.

I then proceeded to arrange a set up for the calibration (initially on the optical bench next to MC2, but now relocated to the SP table, and a cart adjacent to it). It consists of the following:

  • He-Ne laser source
  • Y2 2-inch mirror (AR and HR coated for 532nm) glued onto the PZT and mounted on a machined Newport U100P  - see this elog for details.
  • QPD mounted on a translational stage whose micrometers are calibrated in tenths of an inch (in the plots I have scaled this to mm)
  • A neutral density filter (ND = 2.0) which I added so that the QPD amplifier output did not saturate. I considered using a lens as well to reduce the spot size on the QPD but found that after adding the ND filter, it was reasonably small.
  • High-voltage power supply (on cart)
  • Two SR power supplies (for the PZT driver board and my QPD amplifier
  • SR function generator
  • Laser power source
  • Two oscilloscopes
  • Breadboard holding my QPD amplifier circuit

Having set everything up and having done the coarse alignment using the mirror mount, I proceeded to calibrate the X and Y axes of the QPD using the translational stage. The steps I followed were:

  • Centre spot on QPD using coarse adjustment on the mirror mount: I gauged this by monitoring the X and Y voltage outputs on an oscilloscope, and adjusted things till both these went to zero.
  • Used the tilt knob on the translational stage to roughly decouple the X and Y motion of the QPD.
  • Kept Y-coordinate fixed, took the X-coordinate to close to its maximum value (I gauged this by checking where the voltage stopped changing appreciably for changes in the QPD position.
  • Using this as a starting point, I moved the QPD through its X range, noting voltage output of the X-coordinate (and also the Y) on an oscilloscope.
  • Repeated the procedure for the Y-coordinate.
  • Analysis follows largely what was done in these elogs. I am attaching the script I used to fit an error function to the datapoints, this is something MATLAB should seriously include in cftool (note that it is VERY sensitive to the initial guess. I had to do quite a bit of guessing).

The plots are attached, from which the calibration values cited above are deduced. The linear fits for the orthogonal axis were done using cftool. There is some residual coupling between the X and Y motions of the QPD, but I think this os okay my purposes. 

My next step would be to first tweak the orientation of the PZT in the mount while applying a small excitation to it in order to decouple the pitch and yaw motion as best as possible. Once this is done, I can go ahead and calibrate the angular motion of the PZT in mrad/V.

 

                                                            X-Axis                                                                                                                                Y-axis                                                

QPD-XCalib.pdf  QPD-YCalib.pdf

 

 

Attachment 3: Error_Function_Fitting.zip
  8872   Thu Jul 18 16:30:08 2013 KojiUpdateLSCPRMI+Y arm ALS Sensing Matrices

Hmm. I agree that something was funny.
Let's take the matrix without the arms and confirm the measurement is correct.

  8871   Thu Jul 18 15:55:31 2013 JenneUpdateLSCPRMI+Y arm ALS Sensing Matrices

Last night, I took sensing matrix data at various different offsets for the Yarm.  The sensing matrices I measured were of the PRMI, while the Yarm was (a) Held off resonance, (b) Held at ~50% peak power, and (c) Held on resonance.

The dither lines were clear in the MICH and PRCL spectrum, so I think I'm driving hard enough, but something else seems funny, since clearly the REFL165 I and Q signals were not completely overlapping last night.  If they were, we wouldn't have been able to lock the PRMI using REFL 165 I&Q.

Anyhow, here's the data that was taken.  Data folder is ...../scipts/LSC/SensingMatrix/SensMatData/

Yarm off resonance, SensMat_PRMI_1000cts_580Hz_2013-07-18_012848.dat

SensMatMeas_17July2013_PRMI_YarmOffResonance.png

Yarm at ~50% resonance, SensMat_PRMI_1000cts_580Hz_2013-07-18_013937.dat

SensMatMeas_17July2013_PRMI_YarmMidResonance.png

Yarm on resonance, SensMat_PRMI_1000cts_580Hz_2013-07-18_013619.dat

SensMatMeas_17July2013_PRMI_YarmOnResonance.png

 

  8870   Thu Jul 18 15:34:15 2013 Alex ColeUpdateElectronicsPD Frequency Response Update

 [Eric, Alex]

Our RF Switch arrived today, and we mounted it in rack 1Y1 (1st attachment). 

We connect our input fiber and all of our output fibers to our 1x16 optical splitter (2nd attachment). Note that the 75 meter fiber we are using for the splitter's input is in a very temporary position (3rd attachment - it's the spool).

We successfully turned our laser on and tested the optical splitter by measuring output power at each fiber using our Thorlabs PM20 power meter. Data was taken with the laser running at 67.5 mA and 24 degrees Celsius:

Detector name                  Power

REF DET 192 µW
AS55 146 µW
REFL55 180 µW
REFL11 172 µW
MCREFL 133 µW
REFL33 146 µW
REFL165 180 µW
POP22/POP110 182 µW
POP55 193 µW
POX11 123 µW

 

 

Attachment 1: photo_3_(1).JPG
photo_3_(1).JPG
Attachment 2: photo_1_(4).JPG
photo_1_(4).JPG
Attachment 3: photo_2_(4).JPG
photo_2_(4).JPG
  8869   Thu Jul 18 10:50:54 2013 LisaUpdateLSCPRMI+Y arm ALS success!

Quote:

[Koji, Jenne, Manasa, Annalisa, Rana, Nic]

PRMI locked using 3f signals and Y arm brought to resonance using ALS

Fantastico! :-)

  8868   Thu Jul 18 10:47:21 2013 JamieUpdateLSCPRMI+Y arm ALS success!

AWESOME!  You guys rock.

  8867   Thu Jul 18 02:21:41 2013 KojiUpdateLSCPRMI+Y arm ALS success!

[Koji, Jenne, Manasa, Annalisa, Rana, Nic]

PRMI locked using 3f signals and Y arm brought to resonance using ALS


<<Procedure>>

Preparation:

- After we checked the functionarity of the Yarm ALS, both arms were locked with the IR, and aligned by ASS.

- Disengaged the LSC feedback. Approximately aligned the PRM.

- Recorded the current alignment biases. Turned off all of the oplevs.

- Went into the lab, aligned all of the oplevs on the QPDs (except for the SRM).

- Check the locking of the PRMI.

- Once it is locked, go into the lab again and align the POP QPD.

- Check everything of the PRMI LSC/ASC works.

- Misalign PRM by 0.2

- Lock the arm again. Run ASS again.

- Miaslign ETMX.

ALS:

- Lock the Xarm with green. Adjust the beat freq between 30-50MHz.

- Reset Phase Tracker history.

- Check if there is any offset for the ALS. If there is, adjust it to zero.

- Stabilize the arm with the ALS. We should check the sign of the servo before it is cranked up to the nominal.

- Confirm if the offset FM has LPF (30mHz LPF).

- Run excastep for the ALS offset until we find the TEM00 resonance of the IR

- Record the offset at the resonance.

- Step back by 5 count (=100kHz)

PRMI+ALS:

- Started from the offset of -5.

- Aligned the PRM and the PRMI was locked by REFL165I(x0.8)nadQ(x0.2).

- PRM ASC engaged

- Moved the offset to -4 by ezcastep C1:ALS-OFFSETTER2_OFFSET +0.01,100 -s 0.1

- Moved to -3, -2, -1.5, -1. During the sweep PRCL/MICH gain was tweaked so that the gain is reduced.
  Nominal locking gain was PRCL x+2.5/MICH -30 . During the sweep they were +2.2 / -12
  PRCL FM2/4/5 ON, Later FM3/6 turned on and no problem.

- Moved to -0.9, .... , and finally to 0.


NEXT STEP

- Automation of the PRMI+one arm

- PRMI locking with BS/PRM

- Better sensing matrix

- PRMI+two arms

- Use of the DC signals form the transmission monitors. (High power /low power transmon).

 

 

Attachment 1: Screenshot-Untitled_Window1.png
Screenshot-Untitled_Window1.png
Attachment 2: PRMI_Yarm.pdf
PRMI_Yarm.pdf
  8866   Thu Jul 18 01:10:00 2013 kiwamuUpdateGreen LockingALS Y performance with the new whitening filter

 

Awesome !

  8865   Wed Jul 17 22:51:50 2013 KojiUpdateGreen LockingALS Y performance with the new whitening filter

[Manasa Koji]

Summary:
The new whitening filters improved the out-of-loop ALS stability of the Y arm down to 300Hz (20pm_rms in displacement).


- After modifying the whitening filters, the out-of-loop stability of the arms were tested with the IR PDH signals.

- The X arm showed non-stationarity and it made the ALS servo frequenctly fell out of lock.

- For now we decided to use the Y arm for the PRMI+one arm trial.

- The performance of the ALS was tested with several measurements. (attachment 1)

Cyan: Stability of the beatnote frequency with the MC and the arm freely running. The RMS of the day was ~6MHz.

Blue: Sensing limit of the beat box was tested by giving a signal from Marconi. The same amplitude as the X arm beat was given as the test signal.
This yielded the DC output of ~1200 counts.

Green: Out-of-loop estimation of the beatbox performance. This beat note stability was measured by controlling the arm with the IR PDH signal.
Assuming the PDH signal has better SNR than the beat signal, this gives us the out-of-loop estimation of the stability below 150Hz, which is the
unity gain frequency of the ALS loop.
Above 150Hz the loop does not force this noise to the suspension. Just the noise is injected via a residual control gain (<1).

Black: In-loop evaluation of the ALS loop. This becomes the left over noise for the true stability of the arm (for the IR beam).

Red: The arm was brought to the IR resonance using the ALS offset. The out-of-loop stability was evaluated by the IR PDH signal.
This indeed agreed with the evaluation with the other out-of-loop evaluation above (Green) below 150Hz.


Attachment 2 shows the time series data to show how the arm is brought to the resonance.
1 count of the offset corresponds to ~20kHz. So the arm started from 200kHz away from the resonance
and brought to the middle of the resonance.

(Manasa downloaded the 2k sampled data so that we can use this for presentations.)

Attachment 1: ALS_Y_130717.pdf
ALS_Y_130717.pdf
Attachment 2: ALS_Y2_StripTool.png
ALS_Y2_StripTool.png
  8864   Wed Jul 17 22:49:37 2013 KojiUpdateGreen LockingALS Y whitening filter change

[Koji Annalisa]

We did the same mod of the beatbox for the Y arm too. See
http://nodus.ligo.caltech.edu:8080/40m/8855

  8863   Wed Jul 17 16:15:42 2013 Alex ColeConfigurationElectronicsAS Table Additions

[Eric, Alex]

We decided that the POY Table would be a better home for our REF DET (Newport 1611 FC-AC) than the AS Table. We moved the PD to the POY Table (1st attachment) and routed a fiber from our 1x16 Optical Splitter in the OMC_North rack to the POY Table. REF DET's power supply is now located under the POY table (2nd attachment). We left the fiber described in the previous post on the AS Table.

Afterwards, we hooked a fiber up to our laser module to test it (3rd attachment). The laser was not being distributed, just going to one fiber with a power meter at its end. Everything turns out, but we realized we need to read the power supply's manual before continuing. 

 

 

Attachment 1: photo_1_(3).JPG
photo_1_(3).JPG
Attachment 2: photo_2_(3).JPG
photo_2_(3).JPG
Attachment 3: photo_3.JPG
photo_3.JPG
  8862   Wed Jul 17 11:13:36 2013 Alex ColeConfigurationElectronicsAS Table Additions

[Eric, Alex]

For the RFPD frequency response project, we routed the fiber that will connect our REF DET (on the AS table) to our 1x16 optical splitter (in the OMC_North rack), as pictured. (The new fiber is the main one in the picture, which ends at the right edge near REF DET) Note that we secured the fiber to the table in two places to ensure the fiber would remain immobile and out of other optical paths already in place.

At 2:00 we plan to run fiber from our laser module (in rack 1Y1) to our 1x16 optical splitter (in the OMC_North rack) and measure the power output at one of the splitter's output ports. We plan to keep the output power limited to less than 0.5 mW per optical splitter output.

Attachment 1: photo_(1).JPG
photo_(1).JPG
  8861   Tue Jul 16 19:16:12 2013 ranaUpdateDAQNDS2 Status

I have modified the settings on the router that connects our Martian network to the outside world so that one can access the NDS2 server running on megatron:31200.

To get at the data you point your data getting client (Matlab, ligoDV, DTT, etc.) at our router and the megatron port will be forwarded to you:

131.215.115.189:31200

is what you should point to. Now, it should be possible to run DetChar jobs (e.g. our 40m Summary pages) from the outside on some remote server. You can also grab 40m data on your laptop directly by using matlab or python NDS software.

  8860   Tue Jul 16 18:20:25 2013 JenneUpdateCDSProto-ASC implemented in ASS model

The proto-ASC now includes triggering.  I have updated the hacky temp ASC screen to show the DoF triggering.  I have to go, but when I get back, I'll also expose the filter module triggering.  So, for now we may still need the up/down scripts, but at least the ASC will turn itself off if there is a lockloss.

  8859   Tue Jul 16 17:02:41 2013 Alex ColeConfigurationElectronicsAS Table Additions

 [Eric, Alex]

We added our reference photodetector (Newport 1611, REF DET) to the southern edge of the AS table, as pictured. The detector's power supply is located under the southwest corner of the table, as pictured. We have connected the detector to its power supply, and will connect the detector's fiber input and RF output tomorrow.

EDIT: this is about the RFPD frequency response setup...

Attachment 1: photo_1_(1).JPG
photo_1_(1).JPG
Attachment 2: photo_2_(2).JPG
photo_2_(2).JPG
  8858   Tue Jul 16 15:22:27 2013 manasaUpdateCDSFront ends back up

c1sus, c1ioo, c1iscex and c1iscey were down. Why? I was trying to lock the arm and I found that around this time, several computers stopped working mysteriously. Who was working near the computer racks at this time???

I did an ssh into each of these machines and rebooted them sudo shutdown -r now

But then I forgot / neglected/ didn't know to bring back some of the SLOW Controls computers because I am new here and these computers are OLD. Then Rana told me to bring them back and then I ignored him to my great regret. As it turns out he is very wise indeed as the legends say.

So after awhile I did Burtgooey restore of c1susaux (one of the OLD & SLOW ones) from 03:07 this morning. This brought back the IMC pointing and it locked right away as Rana foretold.

Then, I again ignored his wise and very precious advice much to my future regret and the dismay of us all and the detriment of the scientific enterprise overall.

Later however, I was forced to return to the burtgooey / SLOW controls adventure. But what to restore? Where are these procedures? If only we had some kind of electronics record keeping system or software. Sort of like a book. Made of electronics. Where we could LOG things....

But then, again, Rana came to my rescue by pointing out this wonderful ELOG thing. Huzzah! We all danced around in happiness at this awesome discovery!!

But wait, there was more....not only do we have an ELOG. But we also have a thing called WIKI. It has been copied from the 40m and developed into a thing called Wikipedia for the public domain. Apparently a company called Google is also thinking about copying the ELOG's 'find' button.

When we went to the Wiki, we found a "Computer Restart Procedures" place which was full of all kinds of wonderous advice, but unfortunately none of it helped me in my SLOW Controls quest.

 

Then I went to the /cvs/cds/caltech/target/ area and started to (one by one) inspect all of the targets to see if they were alive. And then I burtgooey'd some of them (c1susaux) ?? And then I thought that I should update our 'Computer Restart Procedures' wiki page and so I am going to do so right now ??

And then I wrote this elog.

 

  8857   Tue Jul 16 14:51:09 2013 gautamConfigurationendtable upgradeAI Board-D000186-Modified notches

 I tried shifting the notch frequencies on the D000186-revision D board given to me by Koji. The existing notches were at ~16 kHz and ~32 kHz. I shifted these to notches at ~64 kHz and ~128 kHz by effecting the following changes (see schematic for component numbering) on Channel 8 of the board-I decided to check things out on one channel before implementing changes en masse:

  • R6 and R7 replaced with 511 ohm smts
  • R8 replaced with 255 ohm smt
  • R14 and R15 replaced with 549 ohm smts
  • R16 replaced with 274 ohm smt

=> New notches should be at 66.3 kHz and 131.7 kHz.

I then measured the frequency response of the modified channel using the SR785, and compared it to the response I had measured before switching out the resistors. The SR785 only goes up to 102 kHz, so I cannot verify the 128 kHz notch at this point. The position of the 64 kHz notch looks alright though. I think I will go ahead and switch out the remaining resistors in the evening.

Note 1: These plots are just raw data from the SR785, I have not tried to do any sort of fitting to poles and zeros. I will do this at some point. 

Note 2: All these smts were taken from Downs. Todd helped me locate the non-standard value resistors. I also got a plastic 25-pin D-sub backshells (the spares are in the rack), with which I have fashioned the required custom ribbon cables (40 pin IDC to 25 pin D-sub with twisted ribbon wire, and a short, 10pin IDC to 10pin IDC with straight ribbon wire).

D000186_Frequency_Response.pdf

  8856   Tue Jul 16 13:48:26 2013 SteveUpdatePEMcranes cleaned

Keven and Steve,

The 3 cranes tested and  wiped off as preparation for upcoming vent.

  8855   Tue Jul 16 10:16:23 2013 KojiUpdateGreen LockingBeatbox XARM whitening modified

The X arm whitening filters of the beatbox were modified.
Now we have about 10 times better floor level above 100Hz and ~3 better at 1Hz.


- The previous whitening was zero@1Hz, pole@10Hz, and the DC gain of the unity.
  When the Marconi signal (~30MHz -25dBm) was given to the beatbox (via ZFL-1000LN),
  the DC output of the beatbox was only 140mV (lame). This corresponded to 220 counts in
  the CDS.
(BTW the signals were calibrated by giving frequency deviation of 1kHz is applied at 125Hz.)

- If you compare the analog measurement of the beatbox output and what we see in the I phase signal,
  you can see that we were completely dominated by the ADC noise (attachment 2, blue and red).

- The new whitening is zero@5.2Hz, pole@159Hz, and the DC gain of 10.

- This improved the sensing noise by factor of ten above 100Hz.

- We are stil llimited by the digitizing noise between 3Hz to 100Hz.
  We need steeper whitening like 2nd order from 1Hz to 100Hz.
(and probably at DC too).
  Now the DC amplitude is about 1.4V (and 2200 counts in the CDS).
  So, it is interesting to see how the sensing limit changes by increasing
  the overall gain by factor of 3, and have (zeros@1Hz & poles@10Hz)^2.

  This can be implemented on a proto-daughter board.

- By the way, the performance below 2Hz is now better than the analog one with the previous whitening.
  This improvement might have come from the replacement of the thick film resisters by thin film resisters.
  (See the circuit diagram)


About the nominal power of the beatbox input.

- Marconi (-20dBm 30MHz) was directly connected to the beatbox. The RF output of -15dBm was observed at the delayline output.
- According to the beatbox schematic, the mixer LO and RF inputs were expected to be -9dBm and -19dBm.
- The nominal mixer LO level is supposed to be 7dBm. Therefore the nominal beatbox input should be -4dBm.

- Assuming 23dB gain of the preamp, the PD output is expected to be -27dBm.

- When the PD out is -27dBm, the RF mon is expected to be -5dBm. This is the level of the RF power expected to be seen in the control room.

- The output of the beatbox was measured as the function of the input to the preamp (before the beatbox input).
  With the nominal gain, we should have observed the amplitude of ~170. And it is now 1700 because of the whitening modification.
 

Attachment 1: Beatbox_mod.pdf
Beatbox_mod.pdf
Attachment 2: ALS_whitening.pdf
ALS_whitening.pdf
Attachment 3: Beatbox_input_dependence.pdf
Beatbox_input_dependence.pdf
  8854   Tue Jul 16 01:17:21 2013 JenneUpdateASCPOP QPD calibration attempt

[Rana, Jenne]

I took POP QPD calibration data with a new method, on Rana's suggestion.  I locked the PRMI, and engaged the ASC servo, and then used awggui (x8) to put dither lines on all of the PRMI-relevant optic's ASCPIT and ASCYAW excitation points.  I then took the transfer function of the suspensions' oplev signals (which are already calibrated into microradians) to the POP_QPD signals (which are in counts).  This way, we know what shaking of any optic does to the axis translation as seen by the POP QPD.  We can also infer (from BS or PRM motion for PR3, and ITMX motion for PR2) what the folding mirrors do to the axis translation.  Note that we'll have to do a bit of matrix math to go from, say, PRM tilt effect to PR3 tilt effect on the axis motion.

The data is saved in /users/jenne/PRCL/July152013_POP_TFs.xml .  There is also a .txt file with the same name, in the same folder, listing the frequencies used by the awg.

I'll analyze and meditate tomorrow, when my brain is not so sleepy.

  8853   Mon Jul 15 17:59:31 2013 gautamConfigurationendtable upgradeDAC at 1Y4- Power Spectrum -6.4kHz bandwidth

Quote:

Those 'peaks' for the oscillations seem ridiculously broad.  I think you should look again, really quickly, with smaller bandwidth at, say, the 2kHz oscillation, to make sure it looks reasonable.

 I did just this, and it looks okay to me:

DAC-6.4_kHz_BW_PSD.pdf

  8852   Mon Jul 15 17:20:43 2013 JenneConfigurationendtable upgradeDAC at 1Y4- Power Spectrum -with the right units

Those 'peaks' for the oscillations seem ridiculously broad.  I think you should look again, really quickly, with smaller bandwidth at, say, the 2kHz oscillation, to make sure it looks reasonable.

  8851   Mon Jul 15 17:16:59 2013 JenneUpdateASCPOP QPD calibration attempt

I tried to retake POP QPD calibration data again today.  The MC was mostly fine, but whenever the PRMI unlocked, both ITM watchdogs would trip.  I'm not sure what was causing this, but the ITM alignment wasn't perfect after this kind of event, so I felt like I was continuously locking and realigning the arms to get the alignment back.   Then, after turning on the ASC and tweaking up the PRM alignment for maximum POP110I signal, I had to recenter the QPD, so none of my previously taken data was useful.  Frustrating.  Also, I had recentered the PRMI-relevant oplevs, but I had these weird locklosses even with nicely centered oplevs.

I have given up for the daytime, and will come back to it if there's a spot in the evening when arm measurements aren't going on.

Here is the data from last week, and the data from today.  The micrometer readings have been calibrated into mm, and I have fit a line to the linear-looking region.  Obviously, for the Pitch calibration, I definitely need to take more data.

ChangeSideMicrometer_July2013_calib2.png

ChangeSideMicrometer_July2013_calib1.png

ChangeVertMicrometer_July2013_calib1.png

  8850   Mon Jul 15 16:51:37 2013 AlexConfiguration Planned AS Table addition

 [Eric, Alex]

We are planning to add our reference PD to the southern third of the AS Table as pictured in the attachment. The power supply will go under the table.

Attachment 1: AS_Table_Ref_PD_Addition.pdf
AS_Table_Ref_PD_Addition.pdf
  8849   Mon Jul 15 16:44:46 2013 AlexUpdateOMCOMC North Safety

 [Eric Alex]

We are planning on testing our laser module soon, so we have added aluminum foil and a safety announcement to the door of OMC North. The safety announcement is as pictured in the attachment.

Attachment 1: photo_2_(1).JPG
photo_2_(1).JPG
  8848   Mon Jul 15 15:54:20 2013 gautamConfigurationendtable upgradeDAC at 1Y4- Power Spectrum -with the right units

Quote:

We need the unit of the voltage power spectrum density to be V/sqrt(Hz).
Otherwise we don't understand anything / any number from the plot.

 I redid the measurement with the appropriate units set on the SR785. Power spectral density plots for no output (top), 500Hz, 1000 counts amplitude sine wave (middle) and 2000Hz, 1000 counts amplitude (bottom) are attached, with the right unit on the Y-axis.

 

DACOffPSD.pdf

 

DAC500PSD.pdf

 

DAC2000PSD.pdf

  8847   Mon Jul 15 14:12:15 2013 SteveUpdateVACRGA scan at day 144

 Pumpdown 75, vacuum normal condition at day 144

 

Attachment 1: pd75mRGA144d.png
pd75mRGA144d.png
Attachment 2: pd75m144d.png
pd75m144d.png
  8846   Mon Jul 15 13:51:17 2013 KojiConfigurationendtable upgradeDAC at 1Y4-Max Output and Power Spectrum

We need the unit of the voltage power spectrum density to be V/sqrt(Hz).
Otherwise we don't understand anything / any number from the plot.

  8845   Mon Jul 15 11:51:18 2013 gautamConfigurationendtable upgradeDAC at 1Y4-Max Output and Power Spectrum

 Summary:

I measured the maximum output of the DAC at 1Y4 as well as its power spectrum. The results are as follows (plots below):

  • Maximum amplitude of differential output: + 10V.
  • Power spectrum has a peak at 64 kHz.

Therefore, the gain of the high-voltage amplification stage on the PZT driver boards do not need to be changed again, as the required output range of 0-100V from the DAC board was realised when the input voltage ranged from -10V to +10 V w.r.t ground. The AI board converts the differential input to a single ended output as required by the driver board.

I will now change some resistors/capacitors on the AI board such that the position of the notches can be moved from 16k and 32k to 64k and 128k.

Procedure:

 Max. amplitude measurement

My previous measurement of the maximum output amplitude of the DAC was flawed as I made the measurement using a single channel of the oscilloscope, which meant that the negative pin of the DAC channel under test was driven to ground. I redid the measurement to avoid this problem. The set up this time was as follows:

  • Positive pin of DAC connected to channel 1 of oscilloscope using break out cable and mini-grabber probe
  • Negative pin of DAC connected to channel 2 of oscilloscope
  • Grounds of channels 1 and 2 connected (I just hooked the mini-grabbers together)
  • Measurement mode on oscilloscope set to channel 1 - channel2
  • Used excitation points set up earlier to output a 3 Hz sine wave with amplitude of 32000 counts from channel 9 of the DAC. 

The trace on the oscilloscope is shown below;

max_amp.JPG

So with reference to ground, the DAC is capable of supplying voltages in the range [-10V 10V]. This next image shows all three traces: positive and negative pins of DAC w.r.t ground, and the difference between the two.

max_amp_all_channels.JPG

 Power spectrum measurement

 

I used the SR785 to make the measurement. The set up was as follows:

  • Positive pin of DAC to A-input of SR560
  • Negative pin of DAC to B-input of SR560
  • A-B output to Channel 1 input A of the SR785
  • SR785 configured to power spectrum measurement

Initially, I output no signal to the DAC, and obtained the following power spectrum. The peak at 65.554 kHz is marked.

DACOffPowerSpec.pdf

I then re-did the measurement with a 200 Hz (left) and 2000 Hz(right), 1000 counts amplitude (I had to change the Ch1 input range on the SR785 from -18dBm to -6dBm) sine wave from channel 9 of the DAC, and obtained the following. The peaks at ~64 kHz are marked.

DACOnPowerSpec.pdf    DAC2kPowerSpectrum.pdf

Now that this peak has been verified, I will work on switching out the appropriate resistors/capacitors on the AI board to move the notches from 16k and 32k to 64k and 128k. 

  8844   Sun Jul 14 18:19:00 2013 AnnalisaUpdateGreen LockingArm cavity scan

Yesterday evening Nic and me were in the lab. The Mode Cleaner was unlocked, but after many attempt we could fix it and we did many scans of the Y arm cavity.

Today I was not able to keep the MC locked. Koji helped me remotely, and eventually the MC locked back, but after half an hour of measurements I had to stop.

I made some more scan of the Y arm though. I also tried to do the same for the X arm, but the MC unlocked before the measurement was finished. I'll try to come back in the night.

  8843   Sun Jul 14 17:47:28 2013 KojiUpdateIOOMC WFS maintenance

Annalisa notified me that the MC autolocker could not keep the MC locked.

I found the initial alignment was not good and the MC was too much excited when the WFS kicked in.

There might have been the WFS offset issue due to the miscentering of the spots on the WFS diodes.

I used the usual procedure of the maintenance and it looked OK if I followed the switching procedure the mc autolocker suppoed to do.
http://nodus.ligo.caltech.edu:8080/40m/7452

I still could not get the autolocker running smoothly. I opened mcup script and compared what was the difference
between my manual sequence and what the script did. The only difference was the lines related to MCL.
It was still turning on the filter module. I checked the MCL path and found that the gain was not zero but 1.0.
So now the MCL gain is set to zero. This solved all the remaining issue.

  8842   Sat Jul 13 03:27:20 2013 AnnalisaUpdateGreen LockingY arm caity scan

I started doing a scan of the Y arm cavity with IR with ALS enabled.

ALS servo tuning:

The servo tuning procedure is basically the same as described in elog 8831.

This time I had a stronger beat note(-14 dBm instead of -24 dBm of the last measurement) thanks to a better alignment.

Plot1 shows the Power spectrum of the BEATY_PHASE_OUT. The RMS is smaller by a factor of 2 (400Hz), corresponding to a residual motion of about 25 pm.

Offset setting avity scan

In order to give an offset linearly growing in time, I used the ezcastep script instead of giving the offset in OFFSETTER2. If the ramp time is long enough, it is not necessary to enable the 30mHz filter.

To span 2 FSR, I started from an offset of 450 and I gave a maximum value of 1600 with a delay of 0.2s between two consecutive steps.

Cavity scan

I did a first scan with the cavity well aligned, basically to know the position of the 00 peaks and choose the best offset range (Plot2)

Then I misaligned the TT2, first in PITCH and yhen in YAW, in order to enhance the HOMs. (Plot3 and Plot4)

More investigation and measurements needed. 

 

 

Attachment 1: PowerSpectrum0712
Attachment 2: ScanCavityAligned.pdf
ScanCavityAligned.pdf
Attachment 3: ScanCavityMisPitch.pdf
ScanCavityMisPitch.pdf
Attachment 4: ScanCavityMisYaw.pdf
ScanCavityMisYaw.pdf
  8841   Fri Jul 12 23:13:32 2013 manasaUpdateGreen LockingALS sensor noise

[Annalisa, Koji, Manasa]

In order to improve the ALS stability we went ahead to check if we are limited by the sensor noise of ALS.

What we did:
RF signals similar to the beatnote were given at the RF inputs of the beatbox.
The frequency of the RF signal was set such that I_OUT was zero (zero-crossing point of the beatbox).
We measured the noise spectrum of the phase tracker output.

Measurements:

Plot 1: X ALS noise spectrum
Plot 2: Y ALS noise spectrum

Discussion:

The X arm ALS noise is not limited by the sensor noise...which means we shoudl come up with clever ideas to hunt for other noise sources.
But this does not seem to be the case for the Y arm ALS. The Y arm part of the beatbox is noisy for frequencies < 100Hz.


After looking into the details and comparing the X and Y arm parts of beatbox, it looks that amplitude of the beat signal seem to affect the Y arm ALS noise significantly and changes the noise spectrum.

To do:
Investigate the effect/limitations of amplitude of the beatnote on the X arm and Y arm beatbox.

Attachment 1: X_ALS_0712.pdf
X_ALS_0712.pdf
Attachment 2: Y_ALS_0712.pdf
Y_ALS_0712.pdf
  8840   Fri Jul 12 21:23:42 2013 JenneUpdateASCPOP QPD calibration attempt

These are the data, one plot for when the vertical QPD position was changed, and one for when the horizontal (yaw) QPD position was changed. 

The micrometer is in inches, so 1 unit is 0.1 inches, I believe.

Clearly, I need to redo the measurement and take more data in the linear region.

Attachment 1: ChangeVertMicrometer_July2013.png
ChangeVertMicrometer_July2013.png
Attachment 2: ChangeSideMicrometer_July2013.png
ChangeSideMicrometer_July2013.png
  8839   Fri Jul 12 18:30:20 2013 CharlesUpdateISSRMS Noise from PMC Transmission

Quote:

It would be better to measure the power spectrum density of the fluctuation.
The RMS does not tell enough information how the servo should be.
In deed, the power spctrum density gives you how much the RMS is in the entire or a specific frequency range.

I wanted the RMS noise simply to establish a very rough estimate of thresholds on RMS detectors that will be part of my device. If you refer to elog 8830, I explain it there. Essentially, when the ISS is first engaged, only one of the 2 or 3 filter stages will be active. Internal RMS threshold detection serves to create a logic input to switch subsequent filters to their 'on' stage.

ELOG V3.1.3-